OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -1/3x + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

fine the equation of the line that passes through (-1, 3) and is perpendicular to y = -1/3x + 7. Is that your question?

OpenStudy (anonymous):

yea ..

OpenStudy (anonymous):

Need to find slope of y = -1/3x + 7 so we can determine the slope of the perpendicular line. We also note that two lines are perpendicular when the product of their slopes is equal to -1. To determine the slope, let's put the equation in slope-intercept form given as y=mx+b where m is the slope and b is the y-intercept when x=0 or at point(0,b). slope m = -1/3, b = +7 therefore the slope of the perpendicular line is 3 because (-1/3)(3) = -1 So far, the equation of the perpendicular line is: y = 3x + b now just plug n the point give to find the value of "b" 3 = 3 + b b = 0 the equation then is: y = 3x + 0 y = 3x